OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8)

OpenStudy (amistre64):

subtract each root by x, and provide a scalar "s" to adjust the results with y = s (-10-x) (-6-x) ; such that when x=-8, y=-8

OpenStudy (anonymous):

what?

OpenStudy (amistre64):

.... see post above :/

OpenStudy (anonymous):

i saw that post above

OpenStudy (amistre64):

then youll have to be more concise as to what you are confused about

OpenStudy (amistre64):

subtract each root by x (-10-x) (-6-x) ; what does this expand into?

OpenStudy (anonymous):

-10x and -6x right

OpenStudy (anonymous):

i think i have no idea how to do this at all

OpenStudy (amistre64):

postives, but yes x^2 + 16x +60

OpenStudy (amistre64):

now, when x=-8, we need to know what this produces, so taht we can know what to scale it by to get -8 back out

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

x^2 + 16x +60; when x=-8 would be?

OpenStudy (amistre64):

might be easier to work out in the factored form (-10--8) (-6--8) (-2) (2) = -4

OpenStudy (anonymous):

-8^2 + 16(8) +60 and i have no clue what you just did there

OpenStudy (anonymous):

is what i did correct without the simplifying part

OpenStudy (amistre64):

i inserted -8 into the factored form i started with, just makes life a little simpler :)

OpenStudy (amistre64):

yes, the math will work it the same

OpenStudy (anonymous):

so it is -64+(-128)+60 is that correct

OpenStudy (amistre64):

almost, remember to include all of x into the squared part like this (-8)^2 = 64 +16(-8) = -128 +60 = + 60

OpenStudy (anonymous):

so you add the 60 to the -128 correct

OpenStudy (amistre64):

yes, then add the 64 to finish it off

OpenStudy (anonymous):

so that would be -4

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

good; now, we want a scalar, a multiplier, such that -4 times ? = -8 -4s = -8; what is s?

OpenStudy (anonymous):

wait i got another question it says to Enter your answer as a function beginning with f(x). and -4 correcct

OpenStudy (anonymous):

*correct

OpenStudy (amistre64):

not yet, lets determine the equation first, then we will worry about syntax

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer to your question is -4 right?

OpenStudy (amistre64):

as is: x^2 +16x + 60 gives us -4 when x=-8, but we dont want a -4 ... we want a -8; so the only thing we can do at this point is find a scalar to modify our equation with such that -4s = -8 this will give us the equation that we want sx^2 +16sx +60s so what is "s"?

OpenStudy (anonymous):

wait how do you find out s though

OpenStudy (amistre64):

by solving this: -4s = -8 -4 times "what" is equal to -8

OpenStudy (anonymous):

2

OpenStudy (amistre64):

can you tell me why we need to do that?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

here is why: we are asked What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8) we found the root parts: x^2 + 16x + 60 satisfies the roots; but it does not satisfy the vertex when x=-8, we want y=-8; at the moment we got x=-8, y=-4; so we need to find a scalar to multiply out setup by to get us to the required y=-8 s = 2 is what we need since -4*2 = -8 therefore our function is: 2x^2 +32x +120

OpenStudy (anonymous):

what how did you get the function thing

OpenStudy (anonymous):

is that the answer right there? if so this is the example answer Format using this sample answer:f(x)=(3)(x^2-21x-19) of how it is supposed to be so would it be 2(x^2+32x+120)?

OpenStudy (amistre64):

... we just worked it from start to finish :/ subtract each root by x, and provide a scalar "s" to adjust the results with y = s (-10-x) (-6-x) , s=2 y = 2 (-10-x) (-6-x) = 2(x^2+16x+60) = 2x^2+32x+120

OpenStudy (anonymous):

ohhh ok that makes since so what about my last question

OpenStudy (amistre64):

well, we have 3 equivalent setups to choose from, and it looks like the middle setup matches your required syntax

OpenStudy (amistre64):

f(x) = 2 (x^2+16x+60)

OpenStudy (anonymous):

ok thank you

OpenStudy (amistre64):

good luck ;)